A student pulls a 10 N block along a rough horizontal surface. The horizontal and the vertical components of the force that he applies are 4 N and 3 N, respectively.
What is the normal force?
If the block is in an impending motion, what is the value of the maximum static friction?
What is the coefficient of static friction?
1
Expert's answer
2019-10-28T10:57:25-0400

There is the equilibrium in the vertical direction:


"N+F_y-W=0"

The normal force:


"N=W-F_y=10-3=7 \ N"

The value of the maximum static friction:


"F_s=F_x=4\ N"

The coefficient of static friction:


"\mu_s=\frac{F_s}{N}=\frac{4}{7}=0.57"